OpenStudy (anonymous):

(GIVE FAN AND MEDAL) HELP!!!!!The table below represents the atmospheric temperature at a location as a function of the altitude: Altitude (in thousand feet) x Temperature (in °C) f (x) 15 4 20 −6 25 −16 30 −26

OpenStudy (anonymous):

The average rate of change of the function between x = 15 to x = 25 is ___degrees Celsius per thousand feet and represents the rate of change of temperature per thousand feet.

OpenStudy (anonymous):

@ganeshie8 ????

ganeshie8 (ganeshie8):

"average rate of change" is same as "slope"

ganeshie8 (ganeshie8):

(15, 4) and `(25, -16) find the slope

OpenStudy (anonymous):

WAIT

OpenStudy (anonymous):

-24?

ganeshie8 (ganeshie8):

Easy, slope is given by \[\dfrac{-16-4}{25-15} = \dfrac{-20}{10} = -2\]

OpenStudy (anonymous):

SO, THAT the answer? -2?

ganeshie8 (ganeshie8):

Yep

OpenStudy (anonymous):

thank you!! you think you can help me with more?

ganeshie8 (ganeshie8):

wil try... i need to go for dinner shortly though..

OpenStudy (anonymous):

okay